OpenStudy (anonymous):

∠a and ∠b are supplementary. If m∠a = 2x + 1 and m∠b = 3x - 6, what is the measure of ∠b? HELP PLEASE

OpenStudy (akashdeepdeb):

Do you know what supplementary means? :)

OpenStudy (akashdeepdeb):

@jaycarterlu4 ?

OpenStudy (akashdeepdeb):

Well, since you are not answering, I'll just tell you what it means. If we have two angles let' say A and B and if they are supplementary then they must sum upto 180 degrees. [Also known as linear pair]. In your question you are given ∠a = 2x + 1 and ∠b = 3x - 6 are supplementary. So ∠a+∠b = 180 Now can you calculate it? :)

OpenStudy (anonymous):

so the answer is ?

OpenStudy (anonymous):

Try doing: (2x + 1) + (3x - 6) = 180

OpenStudy (akashdeepdeb):

@jaycarterlu4 I think it'd be better if you calculate that! Did you get the concept? Beacuse that is more important!

OpenStudy (anonymous):

In case you haven't gotten it yet: (2x + 1) + (3x - 6) = 180 2x + 1 + 3x -6 = 180 5x + 5 = 180 5x = 175 x = ?